OpenStudy (anonymous):

how to solve this system of equations using substitution method : 4y=x 3x-y=70

OpenStudy (godlovesme):

K so x =4y right plug 4y in the original eqn(3x-y=70) for x Make sense so far/ :)

jimthompson5910 (jim_thompson5910):

4y = x or x = 4y so we can replace every 'x' in the second equation with 4y 3x-y=70 3(4y)-y=70 ... replace x with 4y 12y - y = 70 solve for y

OpenStudy (anonymous):

thank you so much